- On 8/15/2019, a 3-year forward contract, expiring 8/15/2022, on a non-dividend-paying stock was entered into when the stock price was $55 and the risk-free interest rate was 10.8% per annum with continuous compounding. 1 year later, on 8/15/2020, the stock price becomes $58. What is the "delivery" price of the forward contract entered into on 8/15/2019?

Round your answer to the nearest 2 decimal points. For example, if your answer is $12.345, then enter "12.35" in the answer box.

- Several months ago, XYZ entered into a long forward contract on an asset with no income. XYZ agreed to pay $30 to seller at maturity.

Today, the contract matures in 9 months. The risk-free rate with continuous compounding is 8.5% per annum, the underlying asset price is $38.55.

Calculate the value of the above forward contract.

Round your answer to the nearest 2 decimal points. For example, if your answer is $12.345, then enter "12.35" in the answer box.
Answer #1

the "delivery" price of the forward contract entered into on 8/15/2019 = S0 x erf x t = 55 x e0.108 x 3 = $ 76.05

--------------------------------

the value of the above forward contract = S0 - PV(K) = 38.55 - 30 x e-0.085 x 9/12 = $ 10.40